# Local certification of geometric graph classes

### v4

@@ 1. Introduction
Local certification is an emerging subfield of distributed computing where the goal is to assign short certificates to each of the nodes of a network (some connected graph $G$) such that the nodes can collectively decide whether $G$ satisfies a given property (i.e., whether it belongs to some given graph class $\mathcal{C}$) by only inspecting their unique identifier, their certificate and the certificates of their neighbors. This assignment of certificates is called a proof labeling scheme, and its complexity is the maximum number of bits of a certificate (as a function of the number of vertices of $G$, which is usually denoted by $n$ in the paper). If a graph class $\mathcal{C}$ admits a proof labeling scheme of complexity $f(n)$, we say that $\mathcal{C}$ has local complexity $f(n)$. Proof labelling schemes are distributed analogues of traditional non-deterministic algorithms, and graph classes of logarithmic local complexity can be considered as distributed analogues of classes whose recognition is in NP [7]. The notion of proof labeling scheme was formally introduced by Korman, Kutten and Peleg in [17], but originates in earlier work on self-stabilizing algorithms (see again [7] for the history of local certification and a thorough introduction to the field). While every graph class has local complexity $O(n^{2})$ [17], [^1] Give to every vertex the adjacency matrix of the graph. the work of [13] identified three natural ranges of local complexity for graph classes:
@@
Closer to the topic of the present paper, the authors of [14] obtained proof labeling schemes of complexity $O(\log n)$ for a number of classes of geometric intersection graphs, including interval graphs, chordal graphs, circular-arc graphs, trapezoid graphs, and permutation graphs. It was noted earlier in [15] (which proved various results on interactive proof labeling schemes for geometric graph classes) that the “only” classes of graphs for which large lower bounds on the local complexity are known (for instance non-3-colorability, or some properties involving symmetry) are not hereditary, meaning that they are not closed under taking induced subgraphs. It turns out that an example of a hereditary class with polynomial local complexity had already been identified in [5] a couple of years earlier: triangle-free graphs (the lower bound on the local complexity given there was sublinear). It was speculated in [15] that any class of geometric intersection graphs has small local complexity, as such classes are both hereditary and well-structured.
@@ Results
In this paper we identify a key rigidity property in graph classes and use it to derive a number of linear lower bounds on the local complexity of graph classes defined using geometric or topological properties. These bounds are all best possible, up to $n^{o(1)}$ factors. So for a number of classical hereditary graph classes studied in structural graph theory, topological graph theory, and graph drawing, we show that the local complexity is $\Theta(n)$. These are the first non-trivial examples of hereditary classes (some of our examples are even monotone) with linear local complexity. Interestingly, all the classes we consider are very close to the class of planar graphs (which is known to have local complexity $\Theta(\log n)$ [9, 6]): most of these classes are either subclasses or superclasses of planar graphs. Given the earlier results on graphs of bounded treewidth [12] and planar graphs, it is natural to try to understand which sparse graph classes have (poly)logarithmic local complexity. It would have been tempting to conjecture that any (monotone or hereditary) graph class of bounded expansion (in the sense of Nešetřil and Ossona de Mendez [21]) has polylogarithmic local complexity, but our results show that this is false, even for very simple monotone classes of linear expansion.
We first show that every class of graphs that contains at most $2^{f(n)}$ unlabelled graphs of size $n$ has local complexity $f(n)+O(\log n)$. This implies all the upper bounds we obtain in this paper, as the classes of graphs we consider usually contain $2^{O(n)}$ or $2^{O(n\log n)}$ unlabelled graphs of size $n$.
@@ 2.1. Local certification
The vertices of any $n$-vertex graph $G$ are assumed to be assigned distinct (but otherwise arbitrary) identifiers $(\text{id}(v))_{v\in V(G)}$ from the set $\{1,\ldots,\text{poly}(n)\}$. When we refer to a subgraph $H$ of a graph $G$, we implicitly refer to the corresponding labelled subgraph of $G$. Note that the identifiers of each of the vertices of $G$ can be stored using $O(\log n)$ bits, where $\log$ denotes the binary logarithm. We follow the terminology introduced by Göös and Suomela [13].
@@
Note that our lower bounds hold in the stronger model of locally checkable proofs of Göös and Suomela [13], where in addition the output of $v$ is allowed to depend on $G[v]$, that is $\mathcal{A}(G,P,v)=\mathcal{A}(G[v],P[v],v)$ for any vertex $v$ of $G$.
@@ 2.2. Geometric graph classes
A drawing of a graph $G$ in the plane is a mapping from the vertices of $G$ to distinct points in the plane and from the edges of $G$ to Jordan curves, such that for each edge $uv$ in $G$, the curve associated to $uv$ joins the images of $u$ and $v$ and does not contain the image of any other vertex of $G$. A graph is planar if it has a drawing in the plane with no edge crossings (such a drawing will also be called a planar graph drawing in the remainder). Every planar graph drawing of a graph $G$ gives a clockwise cyclic ordering of the neighbors around each vertex of $G$. We say that that two planar graph drawings of $G$ are equivalent if the corresponding cyclic orderings are the same. A planar graph embedding of a graph $G$ is an equivalence class of planar graph drawings of $G$. Given a planar graph embedding of a graph $G$, all the corresponding (equivalent) planar drawings of $G$ have the same set of faces (but different choices of outerface yield different planar drawings).
@@
Let $R_{1},\ldots,R_{n}$ be the family of unit-disks or the family of unit-squares. Since $G$ is triangle-free and connected we can assume that the $R_{i}$’s are pairwise distinct (since otherwise $G$ consists of a single edge, and the statement certainly holds). Choose one point $v_{i}$ in each region $R_{i}$ that is not included in any region $R_{j}$ for $j\neq i$, and for any two intersecting regions $R_{i}$ and $R_{j}$, choose a point $r_{ij}=r_{ji}$ in $R_{i}\cap R_{j}$. Inside each region $R_{i}$, join $r_{i}$ to all the points $r_{ij}$ by internally disjoint Jordan arcs. We obtain a planar drawing of $G$, and all planar drawings that we can obtain in this way (starting with the family $R_{1},\ldots,R_{n}$) are equivalent: the clockwise cyclic ordering of the neighbors around each vertex is the same in all such drawings (it corresponds to the clockwise ordering of the intersections with other regions along the boundary of a given region). ∎
We will often need to argue that some planar graphs have unique planar embeddings. The following classical result of Whitney will be crucial.
@@ 3. Linear upper bounds for tiny classes
If there is a constant $c>0$ such that for every positive integer $n$, $|\mathcal{C}_{n}|\leq c^{n}$, then the class $\mathcal{C}$ is said to be tiny. This is the case for all proper minor-closed classes (for instance planar graphs), and more generally any class of bounded twin-width (for instance 1-planar graphs). It is also easy to show that for any finitely generated group $\Gamma$ and any finite set of generators $S$, the class of finite subgraphs of $\mathrm{Cay}(\Gamma,S)$ is tiny (this is proved in [2] for induced subgraphs, but the result for subgraphs follows immediately as these graphs have $O(n)$ edges). On the other hand, unit-interval graphs and unit-disk graphs do not form tiny classes as proved in [19].
@@
We now describe the verifier part. Using the spanning tree $T$, the vertices check that they have been given the same value of $n$ and the same word $w$ describing some graph $G^{\prime}\in\mathcal{C}$. The spanning tree $T$ is then used to compute the number of vertices of $G$ and the root $r$ checks that this number coincides with $n$ and the number of vertices of $G^{\prime}$ (this is standard: each vertex $v$ checks that the number of vertices in its rooted subtree $T_{v}$, which was given as a certificate, is equal to the sum of the number of vertices in the rooted subtrees of its children in $T$, plus 1). Then each vertex $v$ verifies that $\pi$ is a local isomorphism from $G$ to $G^{\prime}$, that is, $\pi$ maps bijectively the neighborhood of $v$ in $G$ to the neighborhood of $\pi(v)$ in $G^{\prime}$.
@@
The remainder of the paper consists in proving lower bounds of order $\Omega(n)$ (or $\Omega(n^{1-\delta})$, for any $\delta>0$), for all the classes mentioned in Corollaries 3.2 and 3.3, except the second item (bounded twin-width) and the last item (our quasi-linear lower bound for unit-disk graphs does not apply to triangle-free unit-disk graphs).
@@ 4. Disjointness-expressing graph classes
    the subgraph of $g(L(A),R(B))$ induced by $N_{g(L(A),R(B))}[S]$ is independent444for all $A,A^{\prime},B,B^{\prime}\subseteq\{1,\ldots,N\}$ there is an isomorphism from $g(L(A),R(B))[N_{g(L(A),R(B))}[S]]$ to $g(L(A^{\prime}),R(B^{\prime}))[N_{g(L(A^{\prime}),R(B^{\prime}))}[S]]$ that is the identity on $S$. of the choice of $A$ and $B$ and has at most $s$ vertices; and
@@
Let $(\mathcal{P},\mathcal{A})$ be a proof labeling scheme for the class $\mathcal{C}$ and let $p\colon\mathbb{N}\to\mathbb{N}$ be its complexity. Let $N$ be a positive integer. For every $A\subseteq\{1,\ldots,N\}$, let $\overline{A}=\{1,\dots,N\}\setminus A$ and let $G_{A}=g(L(A),R(\overline{A}))$. Clearly $G_{A}\in\mathcal{C}$ so the verifier $\mathcal{A}$ accepts the proof $P_{A}=P(G_{A})$ on every vertex of $G_{A}$. Let $n$ denote the maximum order of $G_{A}$ for $A\subseteq\{1,\ldots,N\}$.
@@ 5.1. Penny graphs and unit-distance graphs
The construction is described in Figure 2. Let us argue that for $A,B\subseteq\{1,\ldots,N\}$ there is a unique (up to reflection, translation, and rotation) penny representation of $L(A),R(B)$, and $g(L(A),R(B))$ (shown on Figure 2(c)). We can observe that there exists an ordering $\{v_{1},\ldots,v_{r}\}$ of $L(A)$ (resp. $R(B)$) such that $v_{1},v_{2},v_{3}$ is a triangle, and each $v_{i}$ (for $i>3$) has two neighbors $x$ and $y$ in $\{v_{1},\ldots,v_{i-1}\}$ such that $xyv_{i_{0}}$ is a triangle for some $i_{0}<i$. Once we fix the image of the first triangle $v_{1}v_{2}v_{3}$ in the plane (which must form a unit equilateral triangle), there is a unique way to embed $v_{i}$ in the plane: its image must be at distance exactly one of the images of $x$ and $y$. This condition is satisfied by exactly two points in the plane, one of which is already used by $v_{i_{0}}$.
The set $S=\{c_{1},c_{2}\}$ has size 2 and we can observe that the subgraph induced by the neighborhood of $S$ in $g(L(A),R(B))$ is independent from the choice of $A$ and $B$, and has size at most 6. Hence Condition (ii) of disjointness-expressing is satisfied. Moreover $g(L(A),R(B))$ has at most $1218N+$ 12 vertices ensuring Condition (i) with $\kappa=1$. Finally, we can see on Figure 2(c) that $a_{i}$ and $b_{i}$ cannot both exist at the same time since otherwise their images in the plane would coincide. So if there exists $i\in A\cap B$, $a_{i}$ and $b_{i}$ must both exist and $g(L(A),R(B))$ is not a penny graph. On the other hand, if $A\cap B=\emptyset$ then at most one of $a_{i},b_{i}$ exists for every $i$ and $g(L(A),R(B))$ is a penny graph with representation given in Figure 2. So Condition (iii) of disjointness-expressing is satisfied. ∎
@@
The exact same construction in the proof of Theorem 5.1 can be used to show the following. Indeed, conditions (i) and (ii) of the definition of disjointness-expressing trivially hold (as we consider the same graph) and the same rigidity arguments show that $g(L(A),R(B))$ is a unit-distance graph if and only if $A$ and $B$ are disjoint.
@@ 5.2. Subgraphs of the square grid
The construction is described in Figure 3, and is very similar to the one used for penny graphs in Theorem 5.1. We denote by $D^{L}_{i}$ a left-truncated domino containing 4 vertices and 6 edges, attaching to 2 existing vertices on the left, as shown by one red block on the figure. We add such a subgraph $D^{L}_{i}$ in $L(A)$ if and only if $i$ belongs to $A$. Similarly we define a right-truncated domino $D^{R}_{i}$ containing 4 vertices and 6 edges, attaching to 2 existing vertices on the right, as shown by one blue block on the figure, and $D^{R}_{i}$ is present in $R(B)$ if and only if $i$ belongs to $B$. The set $S=\{c_{1},c_{2}\}$ has size 2 and we can observe that the subgraph induced by the neighborhood of $S$ in $g(L(A),R(B))$ is independent from the choice of $A$ and $B$ and has size at most 6, which fulfills Condition (ii) of disjointness-expressing. Moreover the size of $g(L(A),R(B))$ is at most $1820N+$ 18, satisfying Condition (i) with $\kappa=1$. Finally, we claim that if $g(L(A),R(B))$ is a subgraph of the square grid, the blocks $D^{L}_{i}$ and $D^{R}_{i}$ cannot both exist because there is not enough space in the grid to fit two different vertices at their extremities, in a sense that we explain now. Observe that $g(L(A),R(B))$ can be constructed by gluing $C_{4}$’s along their edges. As every edge of the square grid is shared by exactly two $C_{4}$’s, as soon as we embed one $C_{4}$ of $g(L(A),R(B))$ as an induced subgraph of the square grid, there is at most one way to extend this to an embedding of $g(L(A),R(B))$ as an induced subgraph of the square grid. If both $D^{L}_{i}$ and $D^{R}_{i}$ exist in $g(L(A),R(B))$ and this graph is an induced subgraph of the grid, the aforementioned rigidity property implies that two vertices of the grid belong to both of the truncated dominos, which is impossible since these subgraphs are disjoint in $g(L(A),R(B))$. So in this case, $g(L(A),R(B))$ is not a subgraph of the grid. Conversely it is easy to check that when $A$ and $B$ are disjoint, $g(L(A),R(B))$ is indeed an induced subgraph of the grid. This shows Condition (iii). ∎
@@
Figure 4 illustrates the graph used in the proof. That this is the only possible $1$-planar embedding of this graph follows from a result of [18] (about the outer ring of vertices, which is $C_{2N+6}\boxtimes P_{4}$). On the one hand, $L(A)$ has $82N+$ 8 vertices, including the special vertices $c_{1},\ldots,c_{4}$, and the dotted edge $a_{i}a^{\prime}_{i}$ exists if and only if $i\in A$. On the other hand, $R(B)$ has $4310N+$ 43 vertices, including the same four special vertices $c_{1},\ldots,c_{4}$, and the dotted edge $b_{i}b_{i+1}$ exists if and only if $i\in B$. If $A$ and $B$ are disjoint then the graph $g(L(A),R(B))$ is clearly 1-planar. We now prove that the converse also holds. Consider some $i\in A\cap B$ and assume for the sake of contradiction that $g(L(A),R(B))$ is 1-planar. Then the edge $a_{i}a^{\prime}_{i}$ must cross two edges: $b_{i}b_{i+1}$, and one edge incident to the degree-2 common neighbor of $b_{i}$ and $b_{i+1}$, which is a contradiction. Hence this graph is $1$-planar if and only if $A\cap B=\emptyset$. This proves Condition (iii) of disjointness-expressing. Condition (i) is satisfied with $\kappa=1$ because $g(L(A),R(B))$ has order $4712N+$ 47. Finally regarding Condition (ii), by construction the closed neighborhood of $S=\{c_{1},c_{2},c_{3},c_{4}\}$ in $g(L(A),R(B))$ is independent of the choice of $A$ and $B$ and has size $s=20$.
@@
It seems plausible that a generalization of the result of Theorem 5.8 to $k$-planar graphs can be obtained from the same construction by adding for each edge $uv$ a set of $k-1$ new degree-2 vertices adjacent to both $u$ and $v$.

### v5

@@ 1. Introduction
Local certification is an emerging subfield of distributed computing where the goal is to assign short certificates to each of the nodes of a network (some connected graph $G$) such that the nodes can collectively decide whether $G$ satisfies a given property (i.e., whether it belongs to some given graph class $\mathcal{C}$) by only inspecting their unique identifier, their certificate and the certificates of their neighbors. This assignment of certificates is called a proof labeling scheme, and its complexity is the maximum number of bits of a certificate (as a function of the number of vertices of $G$, which is usually denoted by $n$ in the paper). If a graph class $\mathcal{C}$ admits a proof labeling scheme of complexity $f(n)$, we say that $\mathcal{C}$ has local complexity $f(n)$. Proof labelling schemes are distributed analogues of traditional non-deterministic algorithms, and graph classes of logarithmic local complexity can be considered as distributed analogues of classes whose recognition is in NP [10]. The notion of proof labeling scheme was formally introduced by Korman, Kutten and Peleg in [17], but originates in earlier work on self-stabilizing algorithms (see again [10] for the history of local certification and a thorough introduction to the field). While every graph class has local complexity $O(n^{2})$ [17], [^1] Give to every vertex the adjacency matrix of the graph and the list of identifiers of all the vertices. the work of [13] identified three natural ranges of local complexity for graph classes:
@@
Closer to the topic of the present paper, the authors of [14] obtained proof labeling schemes of complexity $O(\log n)$ for a number of classes of geometric intersection graphs, including interval graphs, chordal graphs, circular-arc graphs, trapezoid graphs, and permutation graphs. It was noted earlier in [15] (which proved various results on interactive proof labeling schemes for geometric graph classes) that the “only” classes of graphs for which large lower bounds on the local complexity are known (for instance non-3-colorability, some properties involving symmetry [13] or the diameter [4]) are not hereditary, meaning that they are not closed under taking induced subgraphs. It turns out that an example of a hereditary class with polynomial local complexity had already been identified in [5] a couple of years earlier: triangle-free graphs (the lower bound on the local complexity given there was sublinear). It was speculated in [15] that any class of geometric intersection graphs has small local complexity, as such classes are both hereditary and well-structured.
@@ Results
In this paper we identify a key rigidity property in graph classes and use it to derive a number of linear lower bounds on the local complexity of graph classes defined using geometric or topological properties. These bounds are all best possible, up to $n^{\delta}$ factors, for any $\delta>0$. So for a number of classical hereditary graph classes studied in structural graph theory, topological graph theory, and graph drawing, we show that the local complexity is $\Theta(n)$. These are the first non-trivial examples of hereditary classes (some of our examples are even monotone) with linear local complexity. Interestingly, all the classes we consider are very close to the class of planar graphs (which is known to have local complexity $\Theta(\log n)$ [8, 6]): most of these classes are either subclasses or superclasses of planar graphs. Given the earlier results on graphs of bounded treewidth [12] and planar graphs, it is natural to try to understand which sparse graph classes have (poly)logarithmic local complexity. It would have been tempting to conjecture that any (monotone or hereditary) graph class of bounded expansion (in the sense of Nešetřil and Ossona de Mendez [21]) has polylogarithmic local complexity, but our results show that this is false, even for very simple monotone classes of linear expansion.
We first show that every class of graphs that contains at most $2^{f(n)}$ unlabelled graphs of size $n$ has local complexity $f(n)+O(\log n)$  [^3] We use the standard assumption that the range of unique identifiers of the vertices is $\text{poly}(n)$, see Section 2.1 for more details.. This implies all the upper bounds we obtain in this paper, as the classes of graphs we consider usually contain $2^{O(n)}$ or $2^{O(n\log n)}$ unlabelled graphs of size $n$.
@@ 2.1. Local certification
The vertices of any $n$-vertex graph $G$ are assumed to be assigned distinct (but otherwise arbitrary) identifiers $(\text{id}(v))_{v\in V(G)}$ from the set $\{1,\ldots,\text{poly}(n)\}$. When we refer to a subgraph $H$ of a graph $G$, we implicitly refer to the corresponding labelled subgraph of $G$. Note that the identifiers of each of the vertices of $G$ can be stored using $O(\log n)$ bits. We follow the terminology introduced by Göös and Suomela [13].
@@
Remark. Note that our lower bounds hold in the stronger model of locally checkable proofs of Göös and Suomela [13], where in addition the output of $v$ is allowed to depend on $G[v]$, that is $\mathcal{A}(G,P,v)=\mathcal{A}(G[v],P[v],v)$ for any vertex $v$ of $G$.
@@ 2.2. Geometric graph classes
A drawing of a graph $G$ in the plane is a mapping from the vertices of $G$ to distinct points in the plane and from the edges of $G$ to Jordan curves, such that for each edge $uv$ in $G$, the curve associated to $uv$ joins the images of $u$ and $v$ and does not contain the image of any other vertex of $G$. A graph is planar if it has a drawing in the plane with no edge crossings (such a drawing will also be called a planar graph drawing in the remainder). Every planar graph drawing of a graph $G$ gives a clockwise cyclic ordering of the neighbors around each vertex of $G$. We say that two planar graph drawings of $G$ are equivalent if the corresponding cyclic orderings are the same. A planar graph embedding of a graph $G$ is an equivalence class of planar graph drawings of $G$. Given a planar graph embedding of a graph $G$, all the corresponding (equivalent) planar drawings of $G$ have the same set of faces (but different choices of outerface yield different planar drawings).
@@
Let $R_{1},\ldots,R_{n}$ be the family of unit-disks or the family of unit-squares. Since $G$ is triangle-free and connected we can assume that the $R_{i}$’s are pairwise distinct (since otherwise $G$ consists of a single edge, and the statement certainly holds). Since $G$ is not a single edge and is triangle-free, we can choose one point $v_{i}$ in each region $R_{i}$ that is not included in any region $R_{j}$ for $j\neq i$, and for any two intersecting regions $R_{i}$ and $R_{j}$, choose a point $r_{ij}=r_{ji}$ in $R_{i}\cap R_{j}$. Inside each region $R_{i}$, join $r_{i}$ to all the points $r_{ij}$ by internally disjoint Jordan arcs. We obtain a planar drawing of $G$, and all planar drawings that we can obtain in this way (starting with the family $R_{1},\ldots,R_{n}$) are equivalent: the clockwise cyclic ordering of the neighbors around each vertex is the same in all such drawings (it corresponds to the clockwise ordering of the intersections with other regions along the boundary of a given region). ∎
We will often need to argue that some planar graphs have unique planar embeddings. We say that a graph is 3-connected if it is connected and remains so after the deletion of any two vertices. The following classical result of Whitney will be crucial.
@@ 3. Linear upper bounds for tiny classes
If there is a constant $c>0$ such that for every positive integer $n$, $|\mathcal{C}_{n}|\leq c^{n}$, then the class $\mathcal{C}$ is said to be tiny. This is the case for all proper minor-closed classes (for instance planar graphs), and more generally any class of bounded twin-width (for instance 1-planar graphs) according to [2]. It is also easy to show that for any finitely generated group $\Gamma$ and any finite set of generators $S$, the class of finite subgraphs of $\mathrm{Cay}(\Gamma,S)$ is tiny (this is proved in [2] for induced subgraphs, but the result for subgraphs follows immediately as these graphs have $O(n)$ edges). On the other hand, unit-interval graphs and unit-disk graphs do not form tiny classes as proved in [19].
@@
We now describe the verifier part. Each vertex checks that it has been given the same value of $n$ and the same word $w$ describing some graph $G^{\prime}\in\mathcal{C}$ as its neighbors. The spanning tree $T$ is then used to compute the number of vertices of $G$ and the root $r$ checks that this number coincides with $n$ and the number of vertices of $G^{\prime}$ (this is standard: each vertex $v$ checks that the number of vertices in its rooted subtree $T_{v}$, which was given as a certificate, is equal to the sum of the number of vertices in the rooted subtrees of its children in $T$, plus 1). Then each vertex $v$ verifies that $\pi$ is a local isomorphism from $G$ to $G^{\prime}$, that is, $\pi$ maps bijectively the neighborhood of $v$ in $G$ to the neighborhood of $\pi(v)$ in $G^{\prime}$.
@@
The remainder of the paper consists in proving lower bounds of order $\Omega(n)$ (or $\Omega(n^{1-\delta})$, for any $\delta>0$), for all the classes mentioned in Corollaries 3.2 and 3.3, except the second and last items of Corollary 3.2.
@@ 4. Disjointness-expressing graph classes
    the subgraph of $g(L(A),R(B))$ induced by $N_{g(L(A),R(B))}[S]$ is independent555for all $A,A^{\prime},B,B^{\prime}\subseteq\{1,\ldots,N\}$ there is an isomorphism from $g(L(A),R(B))[N_{g(L(A),R(B))}[S]]$ to $g(L(A^{\prime}),R(B^{\prime}))[N_{g(L(A^{\prime}),R(B^{\prime}))}[S]]$ that is the identity on $S$. of the choice of $A$ and $B$ and has at most $s$ vertices; and
@@
Let $(\mathcal{P},\mathcal{A})$ be a proof labeling scheme for the class $\mathcal{C}$ and let $p\colon\mathbb{N}\to\mathbb{N}$ be a monotone upper-bound on its complexity. Let $N$ be a positive integer. For every $A\subseteq\{1,\ldots,N\}$, let $\overline{A}=\{1,\dots,N\}\setminus A$ and let $G_{A}=g(L(A),R(\overline{A}))$. Clearly $G_{A}\in\mathcal{C}$ so the verifier $\mathcal{A}$ accepts the proof $P_{A}=P(G_{A})$ on every vertex of $G_{A}$. Let $n$ denote the maximum number of vertices of $G_{A}$ for $A\subseteq\{1,\ldots,N\}$.
@@ 5.1. Penny graphs and unit-distance graphs
The construction is described in Figure 2. Let us argue that for $A,B\subseteq\{1,\ldots,N\}$ there is a unique (up to reflection, translation, and rotation) penny representation of $L(A),R(B)$, and $g(L(A),R(B))$ (shown on Figure 2(c)). We can observe that for each of $L(A),R(B)$, and $g(L(A),R(B))$, there exists an ordering $\{v_{1},v_{2},\ldots\}$ of the vertices such that $v_{1},v_{2},v_{3}$ is a triangle, and each $v_{i}$ (for $i>3$) has two neighbors $x$ and $y$ in $\{v_{1},\ldots,v_{i-1}\}$ such that $xyv_{i_{0}}$ is a triangle for some $i_{0}<i$. Once we fix the image of the first triangle $v_{1}v_{2}v_{3}$ in the plane (which must form a unit equilateral triangle), there is a unique way to embed $v_{i}$ in the plane: its image must be at distance exactly one of the images of $x$ and $y$. This condition is satisfied by exactly two points in the plane, one of which is already used by $v_{i_{0}}$.
The set $S=\{c_{1},c_{2}\}$ has size 2 and we can observe that the subgraph induced by the neighborhood of $S$ in $g(L(A),R(B))$ is independent from the choice of $A$ and $B$, and has size at most 6. Hence Condition (ii) of disjointness-expressing is satisfied. Moreover $g(L(A),R(B))$ has at most $18N+12$ vertices ensuring Condition (i) with $\kappa=1$. Finally, we can see on Figure 2(c) that $a_{i}$ and $b_{i}$ cannot both exist at the same time since otherwise their images in the plane would coincide. So if there exists $i\in A\cap B$, $a_{i}$ and $b_{i}$ must both exist and $g(L(A),R(B))$ is not a penny graph. On the other hand, if $A\cap B=\emptyset$ then at most one of $a_{i},b_{i}$ exists for every $i$ and $g(L(A),R(B))$ is a penny graph with representation given in Figure 2. So Condition (iii) of disjointness-expressing is satisfied. ∎
@@
The exact same construction as in the proof of Theorem 5.1 can be used to deal with unit-distance graphs. Indeed, observe that the same rigidity arguments show that $g(L(A),R(B))$ is a unit-distance graph if and only if $A$ and $B$ are disjoint. Hence by the first item of Remark 4.2 we get the following.
@@ 5.2. Subgraphs of the square grid
The construction is described in Figure 3, and is very similar to the one used for penny graphs in Theorem 5.1. We denote by $D^{L}_{i}$ a left-truncated domino containing 4 vertices and 6 edges, attaching to 2 existing vertices on the left, as shown by one red block on the figure. We add such a subgraph $D^{L}_{i}$ in $L(A)$ if and only if $i$ belongs to $A$. Similarly we define a right-truncated domino $D^{R}_{i}$ containing 4 vertices and 6 edges, attaching to 2 existing vertices on the right, as shown by one blue block on the figure, and $D^{R}_{i}$ is present in $R(B)$ if and only if $i$ belongs to $B$. The set $S=\{c_{1},c_{2}\}$ has size 2 and we can observe that the subgraph induced by the neighborhood of $S$ in $g(L(A),R(B))$ is independent from the choice of $A$ and $B$ and has size at most 6, which fulfills Condition (ii) of disjointness-expressing. Moreover the size of $g(L(A),R(B))$ is at most $20N+18$, satisfying Condition (i) with $\kappa=1$. Finally, we claim that if $g(L(A),R(B))$ is a subgraph of the square grid, the blocks $D^{L}_{i}$ and $D^{R}_{i}$ cannot both exist because there is not enough space in the grid to fit two different vertices at their extremities, in a sense that we explain now. Observe that $g(L(A),R(B))$ can be constructed by gluing $C_{4}$’s along their edges. As every edge of the square grid is shared by exactly two $C_{4}$’s, as soon as we embed one $C_{4}$ of $g(L(A),R(B))$ as an induced subgraph of the square grid, there is at most one way to extend this to an embedding of $g(L(A),R(B))$ as an induced subgraph of the square grid. If both $D^{L}_{i}$ and $D^{R}_{i}$ exist in $g(L(A),R(B))$ and this graph is an induced subgraph of the grid, the aforementioned rigidity property implies that two vertices of the grid belong to both of the truncated dominos, which is impossible since these subgraphs are disjoint in $g(L(A),R(B))$. So in this case, $g(L(A),R(B))$ is not a subgraph of the grid. Conversely it is easy to check that when $A$ and $B$ are disjoint, $g(L(A),R(B))$ is indeed an induced subgraph of the grid. This shows Condition (iii). ∎
@@
Figure 4 illustrates the graph used in the proof. The fact that the graph of Figure 4(c) (even without any of the edges $a_{i}a_{i}^{\prime}$ and $b_{j}b_{j}^{\prime}$) has a unique 1-planar embedding follows from a result of [18] (about the outer ring of vertices, which is $C_{2N+6}\boxtimes P_{4}$). Note that all edges except those on the 2-edge paths between consecutive vertices $b_{j},b_{j+1}$ are crossed exactly once: a pigeonhole argument shows that any other drawing inside the region bounded by the outer ring would have an edge crossed twice. This forces each such 2-edge path between $b_{j}$ and $b_{j+1}$ to lie inside the region bounded by $a_{j},b_{j},a_{j}^{\prime},b_{j+1}$. We emphasize that this is not true if we restrict ourselves to the graph of Figure 4(b), where the 2-edge paths might leave the region bounded by $b_{j}$ and $b_{j+1}$ (for this graph, the 1-planar drawing on the figure is not unique).
On the one hand, $L(A)$ has $2N+8$ vertices, including the special vertices $c_{1},\ldots,c_{4}$, and the dotted edge $a_{i}a^{\prime}_{i}$ exists if and only if $i\in A$. On the other hand, $R(B)$ has $10N+43$ vertices, including the same four special vertices $c_{1},\ldots,c_{4}$, and the dotted edge $b_{i}b_{i+1}$ exists if and only if $i\in B$. If $A$ and $B$ are disjoint then the graph $g(L(A),R(B))$ is clearly 1-planar. We now prove that the converse also holds. Consider some $i\in A\cap B$ and assume for the sake of contradiction that $g(L(A),R(B))$ is 1-planar. Then the edge $a_{i}a^{\prime}_{i}$ must cross two edges: $b_{i}b_{i+1}$, and one edge incident to the degree-2 common neighbor of $b_{i}$ and $b_{i+1}$, which is a contradiction. Hence this graph is $1$-planar if and only if $A\cap B=\emptyset$. This proves Condition (iii) of disjointness-expressing. Condition (i) is satisfied with $\kappa=1$ because $g(L(A),R(B))$ has order $12N+47$. Finally regarding Condition (ii), by subgraph induced by the closed neighborhood of $S=\{c_{1},c_{2},c_{3},c_{4}\}$ in $g(L(A),R(B))$ is independent of the choice of $A$ and $B$ and has size $s=20$.
@@
It seems plausible that a generalization of the result of Theorem 5.8 to $k$-planar graphs can be obtained from the same construction by adding for each edge $uv$ a set of k1k-1 new degree-2 vertices adjacent to both $u$ and $v$.